OpenStudy (i_love_my_nieces):

You have a saving account the offers an annual simple interest rate of 5%. You put $350 into the account. Select the graph showing the the total interest you earn after 1, 2, 3, and 4 years. 1. Graph one 2. Graph two 3. Graph three 4. Graph four

OpenStudy (i_love_my_nieces):

Here are the graph.

OpenStudy (i_love_my_nieces):

@SolomonZelman and @Mr.Dimples

OpenStudy (i_love_my_nieces):

@aprilann143

OpenStudy (whpalmer4):

Do you know the formula to compute simple interest?

OpenStudy (i_love_my_nieces):

Nope.

OpenStudy (i_love_my_nieces):

@whpalmer4

OpenStudy (whpalmer4):

Does it ever occur to you to read the notes or text or watch the video again? It's very unlikely you're expected to magically know the formula without any instructional background. However, to save time, I'll just tell you the formula. \[I = Prt\] where \(I\) is the simple interest earned, \(P\) is the principal balance, \(r\) is the interest rate, expressed as a decimal, and \(t\) is the number of periods (years, in this case). After 1 year, you will earn $300*0.05*1 = <you fill in the number>

OpenStudy (i_love_my_nieces):

150 @whpalmer4 sorry I had to help my mom.

OpenStudy (i_love_my_nieces):

@whpalmer4

OpenStudy (whpalmer4):

What does the term "5%" mean to you?

OpenStudy (whpalmer4):

If I said I would give you 5% of $300, how much would you be expecting to receive?

OpenStudy (i_love_my_nieces):

5% of 300

OpenStudy (whpalmer4):

And how much is that?

OpenStudy (whpalmer4):

Hint, you must be very popular at restaurants if you tip 15% and you think 5% of $300 is $150 :-)

OpenStudy (i_love_my_nieces):

Yeah

OpenStudy (i_love_my_nieces):

So I believe it;s chat B.

OpenStudy (i_love_my_nieces):

graph*

OpenStudy (i_love_my_nieces):

@whpalmer4

OpenStudy (whpalmer4):

B is correct. Not sure how you got it if you can't calculate percentages correctly, but... After one year: \[I = 350*0.05*1 = 17.5\] After two years: \[I = 350*0.05*2 = 35\] After three years: \[I = 350*0.05*3 = 52.5\] After four years: \[I = 350*0.05*4 = 70\] B is the graph that shows the points (1,17.5), (2,35), (3,52.5), (4,70)